The length of a rectangle is 3 ft less than twice the width, and the area of the rectangle is 14 ft squared . Find the dimensions of the rectangle.
Answer by amarjeeth123(569)

Let the width be x ft.
Then the length is (2x-3) ft.
Area of the rectangle=x(2x-3)
x(2x-3)=14
2x^2-3x-14=0
2x^2-7x+4x-14=0
x(2x-7)+2(2x-7)=0
(2x-7)(x+2)=0
x=7/2=3.5 feet.
The width is 3.5 feet and the length is 4 feet.